Suppose that the manager of a firm is planning to meet an order of 1000 units of two products X and Y. The manager's problem is to find the combination of two goods that minimize its cost. He has the firm's cost function of two goods estimated as


C = 5X2 + 20 Y2


By using the Lagrangian multiplier method, find the quantity of X and quantity of Y, subject to X + Y = 1000, that minimize the cost of meeting the order.



1
Expert's answer
2022-01-18T09:59:39-0500

"C=5X^2+20Y^2"

X+Y=1000

Langrangian

"L= C+\lambda(1000-X-Y)"

"=" "5X^2+20Y^2+\lambda(1000-X-Y)"

First order Condition

"\frac{\Delta L}{\Delta X}=10X-\lambda=0" ..........(i)


"X=\frac{\lambda}{10}"


"\frac{\Delta L}{\Delta Y}=40Y-\lambda=0" ...........(ii)


"Y=\frac{\lambda}{40}"

"\frac{\Delta L}{\Delta\lambda}=1000-X-Y=0" ........(iii)


X+Y=1000

Using Eqns (i)) and (ii)

"\frac{\lambda}{10}+\frac{\lambda}{40}=1000"


"\frac{(4\lambda+\lambda)}{40}=1000"

"\lambda=\frac{40,000}{5}= 8000"


"X=\frac{8000}{10}= 800"


"Y=\frac{8000}{40}= 200"